OpenStudy (anonymous):

What is the energy of a photon of light having a wavelength of 4.5 x 10-7 meters? (Planck’s constant is 6.626 x 10-34 joule seconds; the speed of light is 2.998 x 108 m/s)

OpenStudy (michele_laino):

hint: the requested energy is given by the subsequent formula: \[\Large \begin{gathered} E = h\nu = \frac{{hc}}{\lambda } = \hfill \\ \hfill \\ = \frac{{6.62 \cdot {{10}^{ - 34}} \times 3 \cdot {{10}^8}}}{{4.5 \cdot {{10}^{ - 7}}}} = ...joules \hfill \\ \end{gathered} \]